# Clustering by Denoising: Latent plug-and-play diffusion for single-cell data

### v1

@@ 5.2 Analysis of Human Fetal Brain Development Datasets
Preprocessing and Denoising   We analyzed 3,495 cells from Nowakowski et al. (2017) and 15,126 cells from Polioudakis et al. (2019), both of which had undergone QC and transformations analogous to our CITE-seq workflow. Because the studies differ in protocols and development of the tissue under consideration, only 17,638 genes were shared; from this intersection we selected the top 5,000 highly variable genes, yielding matrices $X_{\mathrm{now}}\in\mathbb{R}^{3495\times 5000}$ and $X_{\mathrm{pol}}\in\mathbb{R}^{15126\times 5000}$. Cell-type annotations were available only for the Nowakowski et al. (2017) data, so we performed label transfer to the Polioudakis et al. (2019) data using the Seurat v3 procedure (Stuart et al., 2019) over the shared genes. As $X_{\mathrm{now}}$ exhibited a stronger signal, we trained the diffusion prior on a k=15k\!=\!15 latent projection of it (via PCA) and evaluated DICE on 1,000 randomly sampled cells from $X_{\mathrm{pol}}$ to assess denoising and embedding quality. The training and denoising hyperparameters (noise/annealing schedule, number of diffusion denoising steps, and number of Gibbs iterations) matched those used in the CITE-seq example.
Evaluation   We compare UMAP visualizations for the embeddings constructed from the test data using DICE and the embeddings obtained via the PCA transform learned from the training data. Denoising substantially improves biological interpretability, with two salient gains. The canonical excitatory trajectory $\mathrm{RG}\!\to\!\mathrm{IPC}\!\to\!\mathrm{nEN}\!\to\!\mathrm{EN}$ is visually continuous and easy to follow in the DICE embedding, whereas it appears fragmented under PCA. Further, the cluster of inhibitory neurons (IN) is also more compact and clearly separated after denoising, and DICE reveals a distinct branch emerging from IN in the test dataset that is largely masked in the PCA space. Finally, denoising cleanly separates newborn inhibitory neurons (nIN) from mature IN, highlighting developmental heterogeneity within the inhibitory lineage. Because Nowakowski et al. (2017) dataset is a smaller, high-signal reference spanning fewer cell types, while Polioudakis et al. (2019) dataset is a richer atlas, label transfer yields coarse annotations; consequently, global label-matching can penalize meaningful substructure. Indeed, ARI (using the same embedding clustering scheme as CITE-seq experiment) is $0.42$ for PCA versus $0.35$ for DICE, reflecting over-splitting of broad labels into finer, developmentally coherent groups in the test data. A locality-aware metric is more appropriate here: the cell-type LISI (Local Inverse Simpson’s Index (Korsunsky et al., 2019); lower is better for local purity) computed using the same clusters is markedly lower for DICE ($7.27$) than for PCA ($8.25$), indicating that denoising sharpens lineage relationships and resolves biologically relevant subpopulations that PCA tends to blur.
@@ Discussion
We introduced DICE, a latent plug-and-play framework for denoising and extracting meaningful embeddings from high-dimensional observational data with an underlying low-rank structure. Both synthetic and real-world experiments demonstrated its ability to denoise beyond the training distribution and the ability to leverage clean reference data for denoising. We further showed that DICE can quantify uncertainty in cluster assignments and remains robust under model misspecification. Future work includes extending DICE beyond linear low-rank structures and the i.i.d. noise assumption, as well as improving the efficiency of the sampling procedure.
@@ Appendix A Diffusion Training
Across all experiments, we use the same denoising network, TabularDiffusionMLP, which predicts Gaussian noise $\varepsilon_{t,i}$ from the noised sample $\widehat{U}^{r}_{t,i}\in\mathbb{R}^{k}$ at diffusion step $t$. The architecture is parameterized by the number of residual MLP blocks $M$ and the hidden dimension $D$. For synthetic experiments, we use a smaller model with $M=2$ blocks and hidden dimension $D=64$ ($\approx 150{,}000$ trainable parameters), trained for E=2, [^0] E=2{,}000 epochs. For single-cell experiments, we employ a larger model with $M=8$ blocks and hidden dimension $D=512$ ($\approx 35{,}000{,}000$ trainable parameters), trained for E=20, [^0] E=20{,}000 epochs. We use a batch size of $B=4048$ across all models.
@@
We implemented all experiments in Python. Neural network training was performed in PyTorch (Paszke et al., 2019) on GPU-accelerated hardware, using a mix of AWS g4dn.xlarge EC2 instances and a dedicated Ubuntu server with an NVIDIA GeForce GTX 1070 (8GB RAM).
@@
The runtime of our method can be divided into training and inference. Training is dominated by parameter updates for the diffusion model, with speed depending on hardware and chosen model size. Using the CITE-seq setup as an example, training on 9,000 cells with 3,000 dimensions for 10,000 epochs required approximately 11 hours on the EC2 instance.
Inference proceeds via Gibbs sampling, alternating between the prior and likelihood steps. The prior step, which involves sampling from the diffusion model, is the most computationally expensive. As $\rho$ determines the starting timestep of the diffusion model $t$, the duration of the prior sampling is dependent on $\rho$, larger values of $\rho$ result in longer runtime. For example, one Gibbs iteration on 1,000 test cells takes about 30 seconds for $\rho=1$, corresponding to roughly 30 minutes of total runtime for denoising the test set of the CITE-seq dataset with 100 Gibbs iterations.
@@ Appendix D Discussion on the Clustering Metrics
To assess the quality of clustering in the learned embeddings, we employ three complementary metrics: the average cosine silhouette score (Rousseeuw, 1987), the Adjusted Rand Index (ARI) (Hubert & Arabie, 1985), and the cell-type Locally Invariant Simpson Index (cLISI) (Korsunsky et al., 2019). Below, we provide a brief description of each metric.
@@
Appendix E Ablation Studies on the Role of $\rho$
@@ E.1 Spread across runs
Figure 5: UMAP visualizations of 500 runs of DICE-denoised embeddings for the Cluster 2 center under Setup 1 with increasing $\rho\in[0.1,0.5,1.0,5.0,10.0,20.0]$ (left to right, top to bottom). Input observation in red; training data in grey.
@@
To assess the sensitivity of DICE to different values of $\rho$, we perform an ablation study by denoising a fixed input point multiple times and comparing the spread across runs. By fixing the input observation to a single point, we can isolate the effect of $\rho$ on the variability introduced by the stochastic sampling procedure. We use the two-cluster setup (see section 4, Setup 1), and use the same input point for all runs, the center of Cluster 2. We generate 500 independent runs of DICE for $\rho\in[0.1,0.5,1.0,5.0,10.0,20.0]$.
@@ Results
We visualize the generated embeddings under different values of $\rho$ with a UMAP in Figure 5. As $\rho$ increases, the spread of the sampled points grows: for small $\rho$, points remain tightly concentrated around the cluster center, while for large $\rho$, they disperse more broadly. The spread of the sampled points can be interpreted as the posterior uncertainty under different assumed noise levels. When the assumed noise is low (small $\rho$), the posterior is sharply concentrated around the observation, as the likelihood dominates and little deviation from the measurement is plausible. In contrast, for higher noise levels (large $\rho$), the posterior becomes broader, reflecting greater uncertainty: a wider range of latent points is consistent with the observed data. Thus, varying $\rho$ calibrates how strongly the likelihood influences the denoised embedding.
Figure 6: UMAP visualizations of PCA projections (left) and DICE-denoised embeddings for test points restricted to a single cluster, with increasing values of $\rho\in[1.0,8.0,20.0]$ (left to right). As $\rho$ increases, the embeddings gradually shift from being measurement-aligned to prior-aligned. Training data are shown in grey.
Figure 7: UMAP visualizations PCA projections (left) and DICE-denoised embeddings for test points from a cluster not observed during training, shown for increasing $\rho\in[0.5,2.0,4.0]$ (left to right). With larger $\rho$, the unseen cluster progressively gets mapped to the training clusters, reflecting stronger prior alignment. Training data are shown in grey. The UMAP projection is fitted jointly on the combined training and test data.
@@
We plot the PCA projection (left), and generated denoised embeddings for $\rho\in[1.0,8.0,20.0]$ (left to right) in Figure 6. DICE projects all samples into one of the learned clusters, independent of $\rho$. However, $\rho$ controls which of the learned clusters the observations are assigned to. For small $\rho$, samples remain close to their noisy measurements, resulting in all embeddings staying in the original input cluster. For intermediate $\rho$, most points concentrate within one cluster, but a subset diffuses into neighboring clusters. For large $\rho$, the denoising process aligns closely with the prior, resulting in approximately uniform assignment across clusters, independent of the original observation.
@@
We show a UMAP visualization of PCA projection (left), and generated denoised embeddings for $\rho\in[0.5,2.0,4.0]$ in Figure 7. To observe the test cluster, the UMAP is fitted jointly on the combined training and test data. The PCA projection (left) shows that test measurements form a new cluster, which is non-overlapping with the initial training clusters plotted in gray. For small $\rho$, DICE maintains a distinct cluster of points around the novel cluster. As $\rho$ increases, samples gradually disperse into the nearest training clusters, reflecting the influence of the prior. With $\rho=4.0$, almost all points are assigned to clusters from the training set. Notably, even for large $\rho$, the measurements still exert a directional influence on the embedding. Points are not assigned uniformly inside the clusters, but tend to align with parts of the clusters more consistent with the observed input.
@@ E.3 Discussion
Appendix F Data Sources

### v2

@@ 5.2 Analysis of Human Fetal Brain Development Datasets
Preprocessing and Denoising   We analyzed 3,495 cells from Nowakowski et al. (2017) and 15,126 cells from Polioudakis et al. (2019). After performing standard pre-processing on the raw expression data (see Section H for details), we selected 785 highly variable genes shared by both datasets. This yielded matrices $X_{\mathrm{now}}\in\mathbb{R}^{3495\times 785}$ and $X_{\mathrm{pol}}\in\mathbb{R}^{15126\times 785}$. As $X_{\mathrm{now}}$ exhibited a stronger signal, we trained the diffusion prior on a k=15k\!=\!15 latent projection of it (via PCA) and evaluated DICE on 1,000 randomly sampled cells from $X_{\mathrm{pol}}$ to assess denoising and embedding quality. The training and denoising hyperparameters (noise/annealing schedule, number of diffusion denoising steps, and number of Gibbs iterations) matched those used in the CITE-seq example.
Evaluation   We compare UMAP visualizations for the embeddings constructed from the test data using DICE and the embeddings obtained via the PCA transform learned from the training data. Denoising yielded substantially improved biological interpretability, producing more coherent and homogeneous clusters of similar cell types. A notable improvement is found in the embeddings of the cells belonging to the canonical excitatory trajectory $\mathrm{RG}\!\to\!\mathrm{IPC}\!\to\!\mathrm{nEN}\!\to\!\mathrm{EN}.$ This trajectory is visually continuous and easy to follow in the DICE embedding, whereas it appears fragmented and noisy under PCA. Similar to the CITE-seq example, we benchmarked the DICE embeddings against MAGIC, ALRA, $k$-nearest smoothing, and NMF, using the same set of clustering metrics. The results summarized in 2, show that DICE consistently outperforms all competing denoising methods across all four metrics, validating our claim that diffusion-based denoising sharpens lineage relationships and resolves biologically relevant subpopulations better than existing techniques in the literature.
@@ Discussion
We introduced DICE, a latent plug-and-play framework for denoising and extracting meaningful embeddings from high-dimensional observational data with an underlying low-rank structure. Both synthetic and real-world experiments demonstrated its ability to denoise beyond the training distribution and the ability to leverage clean reference data for denoising. We further showed that DICE can quantify uncertainty in cluster assignments and remains robust under model misspecification. Future work includes 1) extending DICE beyond linear low-rank structures and the i.i.d. noise assumption, 2) improving the efficiency of the sampling procedure, 3) incorporating multimodal data and spatial context, and 4) evaluating the quality of the produced embeddings on clinically meaningful downstream tasks.
@@ Appendix A Diffusion Training
Across all experiments, we use the same denoising network, TabularDiffusionMLP, which predicts Gaussian noise $\varepsilon_{t,i}$ from the noised sample $\widehat{U}^{r}_{t,i}\in\mathbb{R}^{k}$ at diffusion step $t$. The architecture is parameterized by the number of residual MLP blocks $M$ and the hidden dimension $D$. For synthetic experiments, we use a smaller model with $M=2$ blocks and hidden dimension $D=64$ ($\approx 150{,}000$ trainable parameters), trained for E=2, [^0] E=2{,}000 epochs. For single-cell experiments, we employ a larger model with $M=8$ blocks and hidden dimension $D=512$ ($\approx 35{,}000{,}000$ trainable parameters), trained for E=10, [^0] E=10{,}000 epochs. We use a batch size of $B=4048$ across all models.
@@
We implemented all experiments in Python. Neural network training was performed in PyTorch (Paszke et al., 2019) on GPU-accelerated hardware, using a mix of AWS g4dn.xlarge EC2 instances and a dedicated Ubuntu server with an NVIDIA RTX PRO 6000 GPU.
@@
The runtime of our method can be divided into training and inference.
Training is dominated by parameter updates for the diffusion model, with speed depending on hardware and chosen model size. We report the execution times for the Ubuntu server with an NVIDIA RTX PRO 6000 GPU. Using the CITE-seq setup, training on 9,000 cells with 3,000 dimensions for 10,000 epochs required approximately 36 minutes.
Inference proceeds via Gibbs sampling, alternating between the prior and likelihood steps. The prior step, which involves sampling from the diffusion model, is the most computationally expensive. As $\rho$ determines the starting timestep of the diffusion model $t$, the duration of the prior sampling is dependent on $\rho$, larger values of $\rho$ result in longer runtime. The of total runtime for denoising the test set of the CITE-seq dataset with 100 Gibbs iterations and $\rho$ ranging from $5.0$ to $0.5$ took 12 minutes.
@@ Appendix D Discussion on the Clustering Metrics
To assess the quality of clustering in the learned embeddings, we employ five complementary metrics: the average cosine silhouette score (Rousseeuw, 1987), the Adjusted Rand Index (ARI) (Hubert & Arabie, 1985), the cell-type Locally Invariant Simpson Index (cLISI) (Korsunsky et al., 2019), the Normalized Mutual Information (NMI) (Strehl & Ghosh, 2002), and the V-measure (Rosenberg & Hirschberg, 2007). Below, we provide a brief description of each metric.
@@
  4. 4.
    Normalized Mutual Information (NMI). Normalized Mutual Information is an information-theoretic measure for computing the agreement between two sets of cluster labels for the same data points. It measures the agreement between two partitions of the data, $C$ and $T$ as follows:
    $\mathrm{NMI}(C,T)=\frac{2\,I(C;T)}{H(C)+H(T)},$
    where $I(C;T)$ is the mutual information between $C$ and $T$, and $H(\cdot)$ denotes the Shannon entropy (see, Cover & Thomas (2006)).
    The NMI ranges from 0 to $1$, with larger values indicating stronger agreement between the inferred clusters and the biological ground-truth labels. It provides a global summary of clustering accuracy and is widely used for benchmarking atlas-level annotation quality.
  5. 5.
    V-measure. The V-measure provides an entropy-based evaluation scheme of two sets of cluster labels. In our case, these sets are the cluster labels obtained by clustering the denoised embeddings and the original cell-type metadata. The V-measure computes two quantities, namely, homogeneity $h$ and completeness $c$. Let $K$ be the true cell labels and $C$ be the predicted clusters. Then
    $h=1-\frac{H(K\mid C)}{H(K)},\quad\mbox{and}\quad c=1-\frac{H(C\mid K)}{H(C)},$
    where $H(\cdot)$ refers to the Shannon entropy of a distribution. Then, the V-measure is defined as
    $\text{V-measure}=\frac{2\times h\times c}{h+c}.$
    Higher values of V-measure indicate that the reconstructed atlas adheres to the cluster structure encoded in the metadata.
Figure A.1: Visual sketch of the DICE denoising procedure. Training stage: (1) The factor loading matrix $\hat{V}$ is estimated on the reference dataset. (2) A diffusion model is trained to capture the underlying biological variation. Inference stage: (3) The initial latent representation $U^{(0)}$ is obtained using $\hat{V}$. Iteratively: (4) Noise is added to $U^{(s)}$ to produce $Z^{(s)}$. (5) $Z^{(s)}$ is denoised using the learned diffusion model, yielding $U^{(s+w)}$.
Appendix E Sensitivity to $\rho$
@@ E.1 Spread across runs
Figure A.2: UMAP visualizations of 500 runs of DICE-denoised embeddings for the Cluster 2 center under Setup 1 with increasing $\rho\in[0.1,0.5,1.0,5.0,10.0,20.0]$ (left to right, top to bottom). Input observation in red; training data in grey.
@@
To assess the sensitivity of DICE to different values of $\rho$, we perform an ablation study by denoising a fixed input point multiple times and comparing the spread across runs. By fixing the input observation to a single point, we can isolate the effect of $\rho$ on the variability introduced by the stochastic sampling procedure. We use the two-cluster setup (see section 4, Setup 1), and use the same input point for all runs, the center of Cluster 2. We generate 500 independent runs of DICE for $\rho\in\{0.1,0.5,1.0,5.0,10.0,20.0\}$.
@@ Results
We visualize the generated embeddings under different values of $\rho$ with a UMAP in Figure A.2. As $\rho$ increases, the spread of the sampled points grows: for small $\rho$, points remain tightly concentrated around the cluster center, while for large $\rho$, they disperse more broadly. The spread of the sampled points can be interpreted as the posterior uncertainty under different assumed noise levels. When the assumed noise is low (small $\rho$), the posterior is sharply concentrated around the observation, as the likelihood dominates and little deviation from the measurement is plausible. In contrast, for higher noise levels (large $\rho$), the posterior becomes broader, reflecting greater uncertainty: a wider range of latent points is consistent with the observed data. Thus, varying $\rho$ calibrates how strongly the likelihood influences the denoised embedding.
Figure A.3: UMAP visualizations of PCA projections (left) and DICE-denoised embeddings for test points restricted to a single cluster, with increasing values of $\rho\in[1.0,8.0,20.0]$ (left to right). As $\rho$ increases, the embeddings gradually shift from being measurement-aligned to prior-aligned. Training data are shown in grey.
Figure A.4: UMAP visualizations PCA projections (left) and DICE-denoised embeddings for test points from a cluster not observed during training, shown for increasing $\rho\in[0.5,2.0,4.0]$ (left to right). With larger $\rho$, the unseen cluster progressively gets mapped to the training clusters, reflecting stronger prior alignment. Training data are shown in grey. The UMAP projection is fitted jointly on the combined training and test data.
@@
We plot the PCA projection (left), and generated denoised embeddings for $\rho\in\{1.0,8.0,20.0\}$ (left to right) in Figure A.3. DICE projects all samples into one of the learned clusters, independent of $\rho$. However, $\rho$ controls which of the learned clusters the observations are assigned to. For small $\rho$, samples remain close to their noisy measurements, resulting in all embeddings staying in the original input cluster. For intermediate $\rho$, most points concentrate within one cluster, but a subset diffuses into neighboring clusters. For large $\rho$, the denoising process aligns closely with the prior, resulting in approximately uniform assignment across clusters, independent of the original observation.
@@
We show a UMAP visualization of PCA projection (left), and generated denoised embeddings for $\rho\in\{0.5,2.0,4.0\}$ in Figure A.4. To observe the test cluster, the UMAP is fitted jointly on the combined training and test data. The PCA projection (left) shows that test measurements form a new cluster, which is non-overlapping with the initial training clusters plotted in gray. For small $\rho$, DICE maintains a distinct cluster of points around the novel cluster. As $\rho$ increases, samples gradually disperse into the nearest training clusters, reflecting the influence of the prior. With $\rho=4.0$, almost all points are assigned to clusters from the training set. Notably, even for large $\rho$, the measurements still exert a directional influence on the embedding. Points are not assigned uniformly inside the clusters, but tend to align with parts of the clusters more consistent with the observed input.
@@ E.3 Discussion
Appendix F Sensitivity to $k$
Figure A.5: Analysis of Adjusted Rand Index(ARI) for different values of $k$, the dimension of the latent space. x-axis is not to scale.
Setup
To assess the sensitivity of the DICE embeddings to the choice of latent dimension $k$ in Eq. (2), we conducted an ablation study on both the CITE-seq and human neocortex datasets using $k\in\{5,10,15,20,25,50,100,200\}$. The quality of the resulting embeddings was evaluated using the Adjusted Rand Index (ARI).
Results
We summarize our findings in Figure A.5. For the CITE-seq dataset, the ARI values of the DICE embeddings increase up to $k=100$ and decreases thereafter. Furthermore, as the latent dimension grows, the benefit of DICE-based denoising over PCA based embeddings becomes more pronounced. This is expected given the large number of distinct cell types in the dataset, which naturally requires a higher-dimensional latent representation. In contrast, for the human neocortex dataset, the ARI values remain relatively stable across different choices of $k$, with DICE outperforming naive PCA-based embeddings for most values. Based on these observations, we recommend selecting $k$ by identifying an elbow in the scree plot of the singular values; choosing a slightly larger $k$ is generally safe, but excessively large values lead to noisier embeddings and increased computational cost with minimal marginal gain.
Appendix G Data Sources
@@
Appendix H Preprocessing Details for the Single Cell Datasets
H.1 CITE-seq Dataset
We adopt the following pre-processing pipeline for the CITE-seq data obtained from scvi-data. A random stratified sampling was used to subsample 10000 cells from the total pool of cells, maintaining the relative proportion of different cell types.
QC filtering
To remove the cells with poor expression quality and sequencing errors, we adopt the standard QC scheme adopted in Seurat Stuart et al. (2019) and scanpy. We exclude cells with fewer than 200 detected genes and filter out genes expressed in fewer than 3 cells. Additionally, to eliminate cells that may have been dying or stressed at the time of sequencing, we remove all cells with more than 15% mitochondrial RNA content.
Normalization and Batch Effect Removal.
To account for differences in sequencing depth across cells, we perform standard library-size normalization: for each cell, we divide the gene counts by the total UMI count and multiply by $10^{4}$. The resulting values are then transformed using $x\mapsto\log(1+x)$ to stabilize variance and reduce the influence of highly expressed genes. We selected the top 3,000 highly variable genes (Seurat v3 criterion), scaled to unit variance with values capped at 10. Batch effects from sequencing lanes were corrected with Harmony (Korsunsky et al., 2019), yielding an expression matrix $X\in\mathbb{R}^{10000\times 3000}$.
H.2 Human Neo-Cortex Datasets
We analyzed 3,495 cells from Nowakowski et al. (2017) and 15,126 cells from Polioudakis et al. (2019). The pre-processed datasets were obtained from the gEAR (Gene Expression Analysis Resource) portal (Orvis et al., 2021). The downloaded data had already undergone standard QC procedures, as described for the CITE-seq dataset, and the expression values were normalized to a fixed library size and transformed using the $\log 1p$ operation.
Because the two studies differ in experimental protocols and developmental stages, only 17,638 genes were shared across both datasets. From these, we selected the top 5,000 highly variable genes (HVGs) within each dataset and retained the intersection of these HVG sets, yielding 785 shared highly variable genes. This resulted in the matrices $X_{\mathrm{now}}\in\mathbb{R}^{3495\times 785}$ and $X_{\mathrm{pol}}\in\mathbb{R}^{15126\times 785}$.
Cell-type annotations were available only for the Nowakowski et al. (2017) dataset, so we transferred these labels to the Polioudakis et al. (2019) cells using the Seurat v3 label transfer procedure (Stuart et al., 2019) based on the shared set of genes.
Appendix I Implementation Details for Benchmarking
To benchmark DICE embeddings against those generated by popular techniques for scRNA-seq analysis, we compared the performance of our method against MAGIC (van Dijk et al., 2018b), ALRA (Linderman et al., 2022b), NMF (Non-negative Matrix Factorization), scVI (Lopez et al., 2018a) and $k$-nearest neighbor based denoising with $k=5,10$ and $15$. We evaluated clustering performance using Seurat’s FindClusters algorithm with the Leiden method (Traag et al., 2019). To select the optimal hyperparameter, we ran the algorithm over a range of resolution values (0.01–2.0) and retained the clustering that achieved the highest ARI. In the following, we summarize the procedures and their implementation inwour analysis.
  • MAGIC implements a Markov affinity based diffusion smoothing of cellular expression profiles.In our benchmarking experiment, we computed PCA scores from the expression matrix and supplied them directly to MAGIC for denoising. We used the official implementation available at https://github.com/KrishnaswamyLab/MAGIC.
  • ALRA applies an adaptively thresholded low-rank matrix approximation to recover denoised gene-expression values. We employed the pyALRA Python implementation (Lanau & Waterfall, 2025) for all experiments. The expression profiles were provided to ALRA and a low rank approximation was computed. We used 50 PC scores of the denoised matrix as our embeddings for the nechmarking.
  • Non-negative Matrix Factorization (NMF) is a commonly used alternative to PCA for denoising and latent factor extraction in scRNA-seq analysis (Kotliar et al., 2019; Carbonetto et al., 2021). We used the NMF implementation from scikit-learn (Pedregosa et al., 2011), selecting the latent dimension to match the number of PCs used for the DICE embeddings.
  • scVI is a variational-autoencoder–based model that denoises raw UMI counts while accounting for library-size variation and batch effects. For the CITE-seq dataset, we trained scVI with a 10-dimensional latent spacewand used the estimated posterior expectations of the latent variables as our embeddings. We could not apply scVI to the human fetal neocortex dataset because the raw UMI counts were not available.
  • $k$-nearest-neighbor smoothing is a classical local-averaging denoising strategy widely used in pipelines such as Seurat v3 (Stuart et al., 2019). We denoised PCA embeddings by replacing each cell’s representation with the average of its $k$ nearest neighbors (Euclidean distance), using $k\in\{5,10,15\}$.